# Autonomous Constellation Fault Monitoring with Inter-satellite Links: A Rigidity-Based Approach

### v1

@@ 2.2 Rank of the Euclidean Distance Matrices with Fault Satellites
In (Knowles and Gao,, 2023), they observed that there are more than three nonzero singular values when the GCEDM is constructed from an EDM with fault satellites. An example is provided in Figure 1. However, mathematical proofs were not provided in their paper. In this section, we provide and prove several properties of the EDM and GCEDM that are corrupted with faults and measurement noises.
@@ 2.3 Distribution of the singular values of the Geometric Centered EDM
where $\lambda_{i}$ is the $i$th singular value of the geometric centered EDM. If the ISR measurements are completely noiseless, we can detect if a fault satellite exists within the graph by checking if $\gamma_{\text{test}}$ of the GCEDMs is not 0, since GCEDMs have rank 3 when no fault exists. However, when the ISR measurements are noisy, the 4th and 5th singular values increase regardless of faults, as shown in Figure 1 and Corollary 2.2.1. Therefore, to detect faults under the presence of faults, we need to set a threshold that can separate the noise and fault. This threshold needs to be set based on the singular value distributions of the noisy (but non-fault) GCEDMs.
The distribution of the singular values of the three different patterns of noisy GCEDMs is shown in Figure 2. As illustrated in Figure 2, the distribution of $\gamma_{\text{test}}$ is affected by the relative geometry of the satellites. When the satellites are spread in 3D space, the distribution of $\gamma_{\text{test}}$ shifts to the right as the noise magnitude increases, as shown in Figure 2(a) and 2(b). In both non-fault and fault cases, the distribution can be well approximated by a gamma distribution. It is worth mentioning that for some geometries, it is not possible to distinguish fault by looking at the singular values because the distribution does not change in the presence of fault, as shown in Figure 2(c). This corresponds to the case where the non-fault satellites are in the same plane. This observation indicates that it is better to have satellites on diverse orbit planes to effectively detect faults.
@@ 3 Redundantly Rigid Graph-Based Fault Detection
3.1 Test Statistic Distribution Prediction Model
As mentioned in Section 2.3, the distribution of the test statistics of the non-fault noisy GCEDM can be approximated as a gamma distribution, where its shape parameters are determined by the relative geometry of the satellites. To detect faults by checking if $\gamma_{\text{test}}$ exceeds a certain threshold, we are particularly interested in the tails of the distribution of the non-fault case. Therefore, our goal is to develop a model $\hat{f}_{\gamma}$ that can predict the parameters of tails of the $\gamma_{\text{test}}$ (e.g., 99.7 percentile value $\gamma_{\text{test},99.7}$), given the ISR noise magnitude $\sigma_{w}$ and the relative geometry.
In this paper, we propose to use the first three singular values $\lambda_{1},\lambda_{2},\lambda_{3}$ of the GCEDM and the first three columns of the orthogonal matrix $\bm{U}$ of the singular value decomposition $\bm{u}_{1},\bm{u}_{2},\bm{u}_{3}\in\mathbb{R}^{n}(n:\text{number of satellitesin the EDM})$ as the input to the model. These parameters remain almost constant under noise, when $\sigma_{w}$ is sufficiently small compared to the ranges. Compared to using the elements of the GCEDM or EDM as the input, this input is invariant to the permutation of the nodes and is more robust to noise in the ranges. This results in an input dimension of $3(n+1)$, where $n$ is the number of satellites in the EDM. For this paper, we only use EDMs with 6 satellites for fault detection, so the input size is 21. An example of a 2D histogram that shows the mapping between the first three singular values and the 99.7 percentile of the test statistics ($\gamma_{\text{test},99.7}$) is shown in Figure 3. For fault detection, we need a model that can map the input to these percentile values, but from Figure 3, we observe that the map is nonlinear.
In this paper, we trained a neural network with 3 fully connected layers (Table 1) for the predictor
$\gamma_{test,99.7}=\hat{f}_{\gamma}(\bm{z}_{test}),\quad\bm{z}_{test}=\begin{bmatrix}\lambda_{1},\lambda_{2},\lambda_{3},\bm{u}_{1}^{\top},\bm{u}_{2}^{\top},\bm{u}_{3}^{\top}\end{bmatrix}\in\mathbb{R}^{21}$ (10)
The training dataset is obtained by simulating the distribution of $\gamma_{\text{test}}$ for 50000 noisy GCEDMs generated from 6 satellite subgraphs in the target constellation. Note that the training data can be generated by simulation because we assume we know the noise level $\sigma_{w}$ and the configuration of the constellation beforehand. The predictor is trained separately for different constellations so that the predictors have good performance for the geometries that appear frequently in each constellation. In future work, we plan to investigate developing a more general predictor, so that it can generate outputs for arbitrary noise magnitudes and percentile values and a wider set of geometries.
(a) 2D histogram for the Mars constellation case
(b) 2D histogram for the ELFO constellation case
Figure 3: A 2D histogram showing the map between the first three singular values of the GCEDM ($\lambda_{1},\lambda_{2},\lambda_{3}$, x-axes) and the 99.7 percentile (y axes). The values were sampled from a set of noisy EDMs obtained from subgraphs that appear in a Mars and ELFO constellation shown later in Figure 6 and Figure 9. In this paper, we train a 3-layer neural network to obtain the predictor $f^{\gamma}$ that predicts the 99.7 percentile of the test statistics ($\gamma_{test,99.7}$) from the first three singular values (and the first three columns of the orthogonal matrix $\bm{U}$) of the GCEDM. The predictor is trained separately for the two cases as they have significantly different geometries.
Table 1: The configuration of the neural network predictor. The predictor takes in the first three singular values and the elements of its corresponding column vectors and outputs the estimated 99.7 percentile value of the test statistics $\gamma_{\text{test}}$.
Layer Input Size Output Size Activation
1 21 128 ReLU
2 128 32 ReLU
3 32 1 -
@@ 3.2 Clique Listing
In order to construct an EDM and GCEDM, we need a set of range measurements between all pairs of $n$ satellites. However, due to the occultation by the planetary body and attitude constraints, the satellites in the constellation are not fully connected with ISRs in most cases. Therefore, the first step of the fault detection algorithm is to find a set of k-clique subgraphs, which are fully connected subgraphs of $k$ satellites. An example of 5-cliques subgraphs is shown in Figure 4.
Various exact algorithms to find all $k$-cliques have been proposed (Li et al.,, 2020). In this paper, we used the Chiba-Nishizeki Algorithm (Albo) due to its simplicity of implementation (Chiba and Nishizeki,, 1985). For each node $v$, Arbo expands the list of $k$-cliques by recursively creating a subgraph induced by $v$’s neighbors. The readers are referred to (Li et al.,, 2020) for the pseudocode and summary of the $k$-clique listing algorithm. Note that the clique finding algorithm does not have to be executed online. Since the availability of the ISRs can be predicted beforehand using a coarse predicted orbit, we can estimate the topology of the ISRs for future time epochs. We can compute the list of $k$-cliques of the predicted topologies for future time epochs (e.g., one orbit), and uplink them to satellites intermittently. If some of these links were actually not available in orbit, we can remove the $k$-cliques that contain the missing links from the list.
@@ 3.3 Online Fault Detection
Let $\mathcal{L}^{G}_{k,t}$ be the list of all $k$-cliques of the satellite network at time step $t$, and $N_{s}$ the total number of satellites. We propose an online fault detection algorithm, which uses the trained predictor $f^{\gamma}$ and $\mathcal{L}^{G}_{k,t}$. The main routine of the fault detection algorithm is shown in Algorithm 1. The overview of the detection algorithm is as follows. First, we judge whether a subgraph ($k$-clique) contains a fault satellite by comparing the predicted and the observed statistics. If the subgraph is identified as a fault, then we look at the elements of $\bm{u}_{4}$ (the 4th column of the orthogonal matrix obtained by SVD) to find which satellite within the subgraph is generating the fault. In particular, we look for the row with the largest magnitude. The satellite corresponding to this row is registered as a “fault candidate,” and a counter will be incremented. The satellite that has been registered as a fault candidate for most of the time is considered as a fault satellite and is removed from the entire graph. This procedure is repeated until there are no more faults. The key difference compared to the greedy-detection algorithm proposed by (Knowles and Gao,, 2024) is that we are using multiple subgraphs (EDMs) to detect faults. In addition, our algorithm has a predictor that can predict the test statistics for various graph geometries, compared to using a fixed threshold.
@@
Clique Size $k$: The size of the clique needs to be 5 or larger to compute the test statistics. To maximize the number of cliques that can be used for fault detection, we can use $k=5$. However, from observation, we found that when we use $k=5$, it is difficult to determine the fault satellite by looking at the index $\text{argmax}_{i}(\bm{u}_{4})_{i}$. Therefore, we propose to use $k=6$ cliques for fault detection. An example is shown in Figure 5.
Detection time interval $DI=|\bm{t}_{fd}|$: This controls the number of timesteps used to identify fault satellites. In general, the true positive rate will increase with larger DI, because the total number of subgraphs increases and the probability that large test statistics will be observed will increase.
Least required number of fault subgraphs $\delta_{nf}$: The algorithm will not detect faults if the total number of subgraphs detected as faults does not exceed this threshold. This is to avoid judging faults from a small number of samples and to reduce the false alarm rate.
@@
                   $\tilde{\mathbf{G}}\leftarrow-\frac{1}{2}\mathbf{J}^{n}\tilde{\mathbf{D}}\mathbf{J}^{n}$
@@
                   $\gamma_{\text{test},99.7}\leftarrow\hat{f}_{\gamma}(\lambda_{1},\lambda_{2},\lambda_{3},\bm{u_{1}},\bm{u_{2}},\bm{u_{3}})$             $\triangleright$ Predict 99.7 percentile of the test statistic
                   if $\gamma_{\text{test}}>\gamma_{\text{test},99.7}$ then
@@
(a) The geometry and topology of the satellite links. The fault satellite 0 and its connected edges are shown in red. The units are normalized by the Mars radius.
(b) The distribution of the test statistic $\gamma_{\text{test}}$ and the largest row of the 4th column of the orthogonal matrix ($\text{argmax}_{i}|(\bm{u}_{4})_{i}|$). When the clique size is 6, we can detect the fault satellite by looking at the largest index of $\bm{u}_{4}$.
Figure 5: The distribution of the test statistic $\gamma_{\text{test}}$ and the row index which has the maximum element of the 4th singular vector $\bm{u}_{4}$. The geometry and topology of the links were taken from a timestep for the Mars constellation in Figure 6. The noise of $\sigma_{w}=1$ m and fault of $\bar{b}=20$ m is added. As shown in Figure 5(b), we can detect faults from singular vectors more robustly, by using 6-cliques compared 5-cliques.
@@ 4.1 Simulation Configuration and Evaluation Metrics
We validate our proposed algorithms for two different constellations around Mars and the Moon. The detailed configurations of the constellations are shown later in each section. For both simulation cases, the orbits are propagated in the two-body propagator without perturbations. The visibility of the links is calculated assuming that the links are visible when it is not occulted by the planetary body.
We run 500 Monte-Carlo analyses with randomly sampled initial time (within one orbital period) and a set of fault satellites. We tested 60 combinations of different detection and simulation parameters: 3 different numbers of fault satellites (1, 2, or 3), 5 different fault magnitudes ($\bar{b}$ = 5, 8, 10, 15, 20 m), and 4 different detection time intervals $(DI=1,2,3,5)$. The timestep between each detection for $DI\geq 2$ is set to 60 seconds. For the other hyper-parameters, we used $k=6,\delta_{nf}=10,\delta_{rf}=0.2$. Note that the seed of the Monte-Carlo simulation is fixed so that the performance for different parameters is compared among the same conditions except the parameters.
We compute the following metrics to evaluate the performance of the detection algorithm
True Positive Rate($\text{True Positive Rate}\ (TPR)$ $=\frac{TP}{TP+FN}$ (11)
Positive Predictive Value($\text{Positive Predictive Value}\ (PPV)$ $=\frac{TP}{TP+FP}$ (13)
[…]
where $TP,FN,FP,TN$ is the total number of true positives (fault satellite detected as fault), false negatives (fault satellite detected as non-fault), false positives (non-fault satellite detected as fault), and true negatives (non-fault satellite detected as non-fault).
4.2 Case 1: Mars Walker Constellation
The first example is the Walker Delta constellation around Mars, with a total of 12 satellites equally spread to 4 different orbital planes. The orbital elements of the constellation are shown in Table 2. The constellation and the links at $t=0$ are shown in Figure 6(a). The number of 6-cliques in the constellation with the number of subgraphs containing satellite 1,2,3 (satellites in the first plane) is shown in Figure 6(b). For all timesteps, each satellite has at least 32 participations within a 6-clique that can be used for fault detection.
Table 2: Orbital elements of the Mars Walker-Delta constellation (Case 1)
$a$ [km] $e$ [] $i$ [deg] $\Omega$ [deg] $\omega$ [deg] $M$ [deg]
1 15850.55 0.0 60 0 0 [0, 120, 240 ]
2 15850.55 0.0 60 90 0 [114.6, 234.6, 354.6]
3 15850.55 0.0 60 180 0 [229.2, 349.2, 109.2]
4 15850.55 0.0 60 -90 0 [343.8, 103.8, 223.8]
The fault detection results compared to the fixed threshold predictor are shown in Figure 7. For the fixed threshold predictor, $\gamma_{\text{test},99.7}$ was fixed to the 99.7 percentile value of the test statistics, which was computed from 50000 non-fault subgraphs that were randomly sampled from the entire simulation window. We observe that compared to the fixed threshold predictor, the machine learning predictor has a higher TPR, especially at lower fault magnitudes. This is because the predicted $\gamma_{test,99.7}$ is less conservative compared to the fixed threshold, and considers more satellites (subgraphs) as fault candidates. However, the neural network predictor also has a higher FPR than the fixed threshold. This is because since the fixed threshold predictor is conservative, it only considers subgraphs of which the singular values and singular vectors of GCEDMs are largely perturbed, which is likely to be fault satellites. If it desired to make the FPR lower for the neural network predictor, we can do so by increasing the hyperparameter $\delta_{nf}$ and $\delta_{rf}$.
The sensitivity of fault magnitude, detection timesteps, and the number of faults to the detection metrics are shown in Figure 8. In general, the true positive rate is higher and the false positive rate is lower if we have a larger false magnitude and a smaller number of fault satellites, as it would be easier to detect fault subgraphs and distinguish which is the fault satellite. Increasing the detection time interval (DI) increases both the true positive rate and the false positive rate. Therefore, detection length can be used as a hyper-parameter to balance the true positive rate and false positive rate.
(a) Orbital planes with satellite position (black points) and links (orange line) at the initial time epoch. The orbital period is half of Mars’s solar day. The high altitude reduces occultation and enables a dense communication topology.
(b) Transition of the number of 6-cliques over one orbital period. Each satellite has at least 32 of 6 cliques that contain itself throughout the orbit.
Figure 6: Case 1: Walker-Delta constellation around Mars with 12 satellites
Figure 7: Case 1 (Mars constellation): The detection results for different fault magnitudes and the detection time interval (DI). The number of faults is fixed to one. The detector using the fixed threshold is shown in a dashed line, while the detector using the trained neural network is shown in a solid line. The neural network predictor has a higher true positive rate at the cost of having a higher false positive rate since it adaptively changes the threshold and considers more satellites as fault candidates. The fixed threshold is conservative since it is based on the singular value distribution over all subgraphs throughout the orbits.
(a) Detection results with fixed threshold
(b) Detection results with neural network predictor
Figure 8: Case 1 (Mars constellation): Detection results (z-axis) shown as a 3D plot for different fault magnitudes (x-axis), number of detection timesteps (y-axis), and number of faults (planes). The true positive rate decreases with a larger number of faults, since when we have multiple fault satellites that have many shared subgraphs, the first round (satellite removal) of the greedy algorithm can remove the majority of the subgraphs containing the second (and third) false satellite as well, making it difficult to detect the remaining fault satellites. Also, a larger number of faults leads to larger false positive rates, as the 4th singular vector of the GCEDM becomes more perturbed.
4.3 Case 2: Elliptical Lunar Frozen Orbit Constellation
@@
The second simulation scenario is the Elliptical Lunar Frozen Orbit (ELFO) constellation with a total of 12 satellites equally spread to 4 different orbital planes. The orbital elements of the constellation are shown in Table 3. The constellation and the links at $t=0$ are shown in Figure 9(a). The number of 6-cliques in the constellation with the number of subgraphs containing satellite 1,2,3 (satellites in the first plane) is shown in Figure 9(b). In this scenario, each satellite experiences a time window where they have zero or few self-containing subgraphs, where faults cannot be detected.

### v2

@@ 3.2 Rank of the Euclidean Distance Matrices with Fault Satellites
In (Knowles and Gao,, 2023), they observed that there are more than three nonzero singular values when the GCEDM is constructed from an EDM with fault satellites. An example is provided in Figure 4. However, mathematical proofs were not provided in their paper. In this section, we provide and prove several properties of the EDM and GCEDM that are corrupted with faults and measurement noises. In the following proofs, we assume $\bar{f}$ is constant for all satellites.
@@ 3.3 Distribution of the Singular Values of the Geometric Centered EDM
where $\lambda_{i}$ is the $i$th singular value of the geometric centered EDM. If the ISR measurements are completely noiseless, we can detect if a fault satellite exists within the graph by checking if $\gamma_{\text{test}}$ of the GCEDMs is not 0, since GCEDMs have rank 3 when no fault exists. However, when the ISR measurements are noisy, the 4th and 5th singular values increase regardless of faults, as shown in Figure 4 and Corollary 3.2.1. Therefore, to detect faults under the presence of faults, we need to set a threshold based on the singular value distributions of the noisy (but non-fault) GCEDMs that can separate the fault from the noise.
The distribution of the singular values of the three different patterns of noisy GCEDMs is shown in Figure 5. As illustrated in Figure 5, the distribution of $\gamma_{\text{test}}$ is affected by the relative geometry of the satellites. When the satellites are spread in 3D space, the distribution of $\gamma_{\text{test}}$ shifts to the right as the noise magnitude increases, as shown in Figure 5(a) and 5(b). In both non-fault and fault cases, the distribution can be well approximated by a gamma distribution. It is worth mentioning that for some geometries, it is not possible to distinguish fault by looking at the singular values because the distribution does not change in the presence of fault, as shown in Figure 5(c). This corresponds to the case where the non-fault satellites are in the same plane. This observation indicates that it is better to have satellites on diverse orbit planes to effectively detect faults.
@@ 4.1 Clique Listing
In order to construct an EDM and GCEDM, we need a set of range measurements between all pairs of $n$ satellites. However, due to the occultation by the planetary body and attitude constraints, the satellites in the constellation are not fully connected with ISRs in most cases. Therefore, the first step of the fault detection algorithm is to find a set of $k$-clique subgraphs, which are fully connected subgraphs of $k$ satellites. For the graph to be 2-vertex redundantly rigid, we need a subgraph with $k\geq 5$. An example of 5-cliques subgraphs is shown in Figure 6.
Various exact algorithms to find all $k$-cliques have been proposed (Li et al.,, 2020). In this paper, we used the Chiba-Nishizeki Algorithm (Albo) due to its simplicity of implementation (Chiba and Nishizeki,, 1985). For each node $v$, Arbo expands the list of $k$-cliques by recursively creating a subgraph induced by $v$’s neighbors. The readers are referred to (Li et al.,, 2020) for the pseudocode and summary of the $k$-clique listing algorithm. Note that the clique finding algorithm does not have to be executed online. Since the availability of the ISRs can be predicted beforehand using a coarse predicted orbit, we can estimate the topology of the ISRs for future time epochs. We can compute the list of $k$-cliques of the predicted topologies for future time epochs (e.g., one orbit), and uplink them to satellites intermittently. If some of these links were actually not available in orbit for some reason, we can remove the $k$-cliques that contain the missing links from the list.
@@ 4.2 Online Fault Detection
Let $\mathcal{L}^{G}_{k,t}$ be the list of all $k$-cliques of the satellite network at time step $t$, and $N_{s}$ the total number of satellites. We propose an online fault detection algorithm, which uses $\mathcal{L}^{G}_{k,t}$. The main routine of the fault detection algorithm is shown in Algorithm 1. The overview of the detection algorithm is as follows. First, we judge whether a subgraph ($k$-clique) contains a fault satellite by comparing the user-set threshold $\bar{\gamma}_{\text{test}}$ and the observed test statistics value. If the subgraph is identified as a fault, then we look at the elements of $\bm{u}_{4}$ (the 4th column of the orthogonal matrix obtained by SVD) to find which satellite within the subgraph is generating the fault. In particular, we look for the row with the largest magnitude. The satellite corresponding to this row is registered as a “fault candidate,” and a counter will be incremented. The satellite that has been registered as a fault candidate for most of the time is considered as a fault satellite and is removed from the entire graph. This procedure is repeated until there are no more faults. The key difference compared to the greedy GPS fault-detection algorithm proposed by (Knowles and Gao,, 2024) is that we are using multiple subgraphs (EDMs) to detect faults, instead of using a (large) single fully-connected graphs.
@@
Clique Size $k$: The size of the clique needs to be 5 or larger to compute the test statistics. The clique size $k=5$ maximizes the number of cliques that can be used for fault detection. However, from observation, we found that when we use $k=5$, it is difficult to determine the fault satellite by looking at the index $\text{argmax}_{i}(\bm{u}_{4})_{i}$. Therefore, we propose to use $k=6$ cliques for fault detection. An example is shown in Figure 7.
Detection time interval $DI=|\bm{t}_{fd}|$: This controls the number of timesteps used to identify fault satellites. In general, the true positive rate (TPR) will increase with larger DI, because the total number of subgraphs increases and the probability that large test statistics will be observed will increase.
Least required number of fault subgraphs $\delta_{nf}$: The algorithm will not detect faults if the total number of subgraphs detected as faults does not exceed this threshold. We set this threshold to avoid judging faults from a small number of samples and to reduce the false alarm rate.
@@
Test statistics threshold $\bar{\gamma}_{\text{test}}$: This controls the trade-off between increasing the TPR and decreasing the FPR. One way to set this threshold is to sample a sufficiently large number of singular values from subgraphs over different epochs, and compute its $x$ percentile value close to the tail (e.g., 99 percentile). When a small $\bar{\gamma}_{\text{test}}$ is used, more subgraphs are listed as fault subgraphs, which increases the probability of fault satellites being detected, but also increases the probability of listing normal satellites as a fault. The threshold $\bar{\gamma}_{\text{test}}$ and the detection time interval $DI$ need to be determined by pair by the system designer to get the desired balance between TPR and FPR.
@@
$\quad\quad\bar{\gamma}_{\text{test}}$                   $\triangleright$  Hyper-parameter: Test statistics threshold
@@
                   $\tilde{\mathbf{G}}\leftarrow-\frac{1}{2}\mathbf{J}^{k}\tilde{\mathbf{D}}\mathbf{J}^{k}$
@@
                   if $\gamma_{\text{test}}>\bar{\gamma}_{\text{test}}$ then
@@
(a) The geometry and topology of the satellite links. The fault satellite 0 and its connected edges are shown in red. The units are normalized by the Moon radius.
(b) The distribution of the test statistic $\gamma_{\text{test}}$ and the largest row of the 4th column of the orthogonal matrix ($\text{argmax}_{i}|(\bm{u}_{4})_{i}|$). When the clique size is 6, we can detect the fault satellite $s_{f}=0$ by counting the number of subgraphs where the largest index of $\bm{u}_{4}$ corresponds to $s_{f}$.
Figure 7: The distribution of the test statistic $\gamma_{\text{test}}$ and the row index which has the maximum element of the 4th singular vector $\bm{u}_{4}$. The geometry and topology of the links were taken from a timestep for the lunar constellation in Figure 8. The noise of $\sigma_{w}=1$ m and fault of $\bar{b}=20$ m is added. As shown in Figure 7(b), we can detect faults from singular vectors more robustly, by using 6-cliques compared 5-cliques.
@@ 5.1 Simulation Configuration and Evaluation Metrics
We validate our proposed algorithms for the constellations around the Moon.
5.1.1 Constellation
We consider an Elliptical Lunar Frozen Orbit (ELFO) constellation with a total of 12 satellites equally spread to 4 different orbital planes. The orbital elements of the constellation are shown in Table 1, and the 3-dimentional plots of the constellation and the available ISls at $t=0$ are shown in Figure 8(a). The orbits are propagated in the two-body propagator without perturbations.
@@ 5.1.1 Constellation
The visibility of the links is calculated assuming that the links are visible when it is not occulted by the Moon. The number of 6-cliques in the constellation with the number of subgraphs containing satellite 1,2,3 (satellites in the first plane) is shown in Figure 8(b). Each satellite experiences a time window where they have zero or few self-containing subgraphs, where faults cannot be detected.
@@
(b) Number of 6-cliques for different time epochs. The satellite near the perilune has few or zero 6-cliques containing itself due to the occultation by the Moon. Sat 1, 2, 3 are the 3 satellites in plane 1.
Figure 8: Elliptical lunar frozen orbit constellation around Moon with 12 satellites.
5.1.2 Simulation Parameters
We assumed that the two-way noise measurement error is $\sigma_{w}=1$ m. We tested 180 combinations of different detection and simulation parameters: 3 different numbers of fault satellites (1, 2, or 3), 5 different fault magnitudes ($\bar{b}$ = 5, 8, 10, 15, 20 m), and 4 different detection time intervals $(DI=1,2,3,5)$, and 3 different test statistic thresholds $\bar{\gamma}_{\text{test}}$. The timestep between each detection for $DI\geq 2$ is set to 60 seconds.
The three thresholds are computed by sampling the test statistics from all the (non-fault but noisy) 6-cliques for one orbital period, and computing the $95,99$, and $99.9$ percentile value. The test statistics are sampled every 60s for one orbital period, which results in total of 256742 subgraphs. The computed thresholds for $95,99$, and $99.9$ percentile are $\bar{\gamma}_{\text{test}}^{95}=3.58\times 10^{-7},\bar{\gamma}_{\text{test}}^{99}=4.57\times 10^{-7},$ and $\bar{\gamma}_{\text{test}}^{99.9}=5.86\times 10^{-7}$, respectively. The distribution and the percentile values of the sampled test statistics are shown in Figure 9.
For each of these cases, we run 500 Monte-Carlo analyses with randomly sampled initial time (within one orbital period) and a set of fault satellites. For the other hyper-parameters, we used $k=6,\delta_{nf}=10,\delta_{rf}=0.2$. Note that the seed of the Monte-Carlo simulation is fixed so that the performance for different parameters is compared among the same conditions except the parameters.
Figure 9: The cumulative distribution of the sampled test statistics $\gamma_{test}$. The test statistics are sampled from all 6-cliques once every 60s for one orbital period. The 95, 99, and 99.9 percentiles are used as thresholds ($\bar{\gamma}_{\text{test}}^{95},\bar{\gamma}_{\text{test}}^{99},\bar{\gamma}_{\text{test}}^{99.9}$) to detect faults.
5.1.3 Evaluation Metrics
We compute the following metrics to evaluate the performance of the fault detection algorithm
True Positive Rate or Recall($\text{True Positive Rate or Recall}\ (TPR)$ $=\frac{TP}{TP+FN}$ (10)
$P_{4}\ \text{metric}\ (P_{4})$ $=\frac{4\cdot TP\cdot TN}{4\cdot TP\cdot TN+(TP+TN)\cdot(FP+FN)}$ (12)
where $TP,FN,FP,TN$ is the total number of true positives (fault satellite classified as fault), false negatives (fault satellite classified as non-fault), false positives (non-fault satellite classified as fault), and true negatives (non-fault satellite classified as non-fault). $P_{4}$ value is the harmonic mean of four key conditional probabilities (precision, recall, specificity, and negative predictive value), and $P_{4}=1$ requires all of these values to be 1 (Sitarz,, 2023).
5.2 Simulation Result